On December 31, 2015, Merlin Company had outstanding 400,000 shares of common stock and 40,000 shares of 8% cumulative preferred stock (par $10). On February 28, 2016, Merlin issued an additional 36,000 shares of common stock. A 10% stock dividend was declared and distributed on July 1, 2016. On September 1, 2016, 9,000 shares were retired. At year-end, there were fully vested incentive stock options outstanding for 30,000 shares of common stock (adjusted for the stock dividend). The exercise price was $18. The market price of the common stock averaged $20 during the year. Also outstanding were $1,000,000 face amount of 10% convertible bonds issued in 2013 and convertible into 50,000 common shares (adjusted for the stock dividend). Net income was $900,000. The tax rate for the year was 40%.

Required: Compute basic and diluted EPS (rounded to 2 decimal places) for the year ended December 31, 2016.

You must show your computations to receive any credit.

Hint:

Basic = $1.85

Diluted = $1.77
Answer #1

Answer:

Basic EPS = ($900,000 - $32,000) / ( 400,000 * 1.10 + 36000*(10/12)*1.10 - 9000*(4/12)  
= $868,000/470,000
= $1.85
Diluted EPS = ($900,000 - $32,000 +$100,000 -40% *($100,000)) / (400000*1.10 + 36000*(10/12)*1.10 - 9000*(4/12) + (30000-27000) + 50000)
= $928,000/ 523,000 = $1.77
Shares reaquired for diluted EPS
(30,000 x $18)/ 20 = $540,000 / 20 = 27,000 shares